i need to know:
when the problem is log base2 X + log base2 (X + 7) = 3
doesnt that need to be rewritten and then solved?

Answer by venugopalramana(3286)

i need to know:
when the problem is log base2 X + log base2 (X + 7) = 3
doesnt that need to be rewritten and then solved?
LOG[X(X+7)] TO BASE 2 =3
X(X+7)=2^3=8
X^2+7X-8=0
X^2+8X-X-8=0
X(X+8)-1(X+8)=0
(X-1)(X+8)=0
X=1..OR.......-8.........
SING LOG OF NEGATIVE NUMBER DOES NOT EXIST
X=1 IS THE ANSWER